OpenStudy (anonymous):

What is the solution to the equation 2 ^1 – 2x = 3 ?

OpenStudy (ash2326):

@melissaneedshelp is this your question \[2^{1-2x}=3\] ?

OpenStudy (anonymous):

yes!

OpenStudy (ash2326):

@melissaneedshelp Have you studied logarithms?

OpenStudy (anonymous):

roughly

OpenStudy (ash2326):

Apply \(\log_2\) to both the sides!!!

OpenStudy (ash2326):

And tell me what do you get, If you need any help do let me know

OpenStudy (anonymous):

\[^{1-2x}= 1??\]

OpenStudy (ash2326):

Oh you made a little mistake We have \[2^{1-2x}= 3\] Apply\(\log_2\) to both the sides \[\log_2 {2^{1-2x}}= \log_2 3\] Now we know \[\log_a a^b=b\] so \[1-2x=\log_2 3\] Now we need to look into logarithm tables or use calculator to find \(\log_2 3\) Can you solve now?